Journalist Report – January 14th

English

Title: Los caminos de la vida (Colombian song)

A hard day begins, two EVAs, a Greenhab that needs to be reborn, a suit that needs repair, the cleaning of the habitat and the dome science. Added to this, the cold Martian begins to make that some of our crew from the tropic land are attacked by strong colds, however, it seems that neither these harsh days nor Martian hostility managed to silence the laughter that comes from spontaneous comments.

First EVA of the day, three of our crew members accompanied by MARStha Leticia, go out to explore and collect material to start their science projects, meanwhile the other 3 crew stay in the habitat carrying out maintenance activities that do not wait. After this EVA again we all meet to taste the culinary creations of our commander and quickly we are ready to change roles.

At the end of the last EVA the tiredness inevitably becomes noticeable, we begin to value the earthly comforts, but at the same time we continue to marvel at the inspiring beauty of our current home. We know that we do not arrive at Mars by chance and that we are sharing this small space of the universe with people with the same dream, reaching the stars.

With little modesty many Colombians often say they are part of the happiest country on Earth, a statement that becomes controversial when we enter to discuss the definition of happiness, what is indisputable is that our new Colombo-Martians have achieved in a short time to cheer Cold Martian desert and fitting to any definition of happiness make it one of the happiest places in the Solar System.

Spanish

Titulo: Los caminos de la vida

Un día duro empieza, dos EVAs, un Greenhab que necesita renacer, un traje que necesita reparación, la limpieza del hábitat y del domo ciencia. Sumado a esto, el frio marciano empieza a hacer de las suyas y algunos de nuestros tripulantes provenientes del trópico terrestre se ven atacados por fuertes resfriados, sin embargo, parece que ni estas duras jornadas ni la hostilidad marciana lograran silenciar las carcajadas que nacen de comentarios espontáneos.

Primer EVA del día, tres de nuestros tripulantes acompañados de MARStha Leticia, salen a explorar y a recoger material para iniciar sus proyectos de ciencia, mientras tanto los otros 3 tripulantes se quedan en el habitad adelantando actividades de mantenimiento que no se hacen esperar. Finalizado este EVA nuevamente todos nos reunimos a degustar las creaciones culinarias de nuestro comandante y rápidamente nos disponemos a cambiar de roles.

Al terminar el ultimo EVA el cansancio inevitablemente se hace notar, empezamos a valorar las comodidades terrícolas, pero al mismo tiempo seguimos maravillando de la belleza inspiradora de nuestro hogar actual. Sabemos que no llegamos a Marte por casualidad y que estamos compartiendo este pequeño espacio del universo junto a personas con un mismo sueño, el de alcanzar las estrellas.

Con poca modestia muchos colombianos suelen decir que hacen parte del país más feliz de la Tierra, una afirmación que se vuelve polémica cuando entramos a discutir la definición de felicidad, lo que es indiscutible es que nuestros nuevos Colombo-marcianos han logrado en poco tiempo alegrar frio desierto marciano y encajándose a cualquier definición de felicidad hacerlo uno de los lugares más felices del Sistema Solar.

EVA Report – January 14th

EVA # 1

Author: Liza Forero

Purpose of EVA:
Recognition of the work area, practice in the use of suits and rovers and implementation of the protocol for taking samples.

Start time:
11:00 am

End time:
1:04 pm

Narrative:
We divided the EVA of the day in two groups, each one of 3 people. Each EVA has its own mission expert and a mission specialist. The mission expert for EVA #1 is Yael Méndez and the mission expert is Oscar Ojeda.

Samples were taken from two different areas for microbiological purposes, two from Phobos Peak and the other two were taken besides the 1103 road.

Before the EVA end, three samples of regolith were taken from Overlook Ridge to begin with the germination viability tests.

Destination: Phobos Peak.

Coordinates:
Sample coordinates:

01142019 – M1: 518419, 4250481
01142019 – M2: 518412, 4250427
01142019 – M3: 518956, 4251197
01142019 – M4: 518987, 4251193

Participants:
Yael Méndez, Oscar Ojeda, Hermes Bolívar

Road(s) and routes per MDRS Map:
For both zones we took the 1103 road and to arrive to the exact point of the samples for Phobos Peak we walk NE for 4 to 5 minutes.

Mode of travel: Rover and walking

EVA # 2

Author: Liza Forero

Purpose of EVA:
Recognition of the work area, practice in the use of suits and rovers and implementation of the protocol for taking samples.

Start time:
2:00 pm

End time:
3:56 pm

Narrative:
The mission expert for EVA #2 is Liza Forero and the mission specialist is David Mateus.

Samples were taken from two different areas for microbiological purposes. Three samples from Phobos Peak and one sample from the URC North Site.

Before the EVA end, three samples of regolith were taken from Overlook Ridge to begin with the germination viability tests.

Destination: Phobos Peak and URC North Site.

Coordinates:
Sample Coordinates:

01142019 – M5: 518530, 4250496
01142019 – M6: 518540, 4250475
01142019 – M7: 518654,4250439
01142019 – M8: 518936, 4251177

Participants:
Liza Forero, David Meteus, Fredy Castaneda

Road(s) and routes per MDRS Map:
For both zones we took the 1103 road and to arrive to the exact point of the samples for Phobos Peak we walk NE for 5 minutes.

Mode of travel:
Rovers and walking.

GreenHab Report – January 14th

[category GreenHab-report]

Greenhab Officer:

Hermes Bolivar

Environmental Control:

Heating

Cooling w/ ambient air (3 hrs)

40% Shade Cloth on

80% Shade Cloth on

Average Temperatures:

Low: 16.6°C

High: 31.3°C

Hours of Supplemental Light:

5

Daily Water Usage of Crops:

10 gallons

Water in Blue Tank (gallons):

87.3%, 262 Gallons.

Times of Watering for Crops:

9:12

11:40

15:03

17:04

19:02

Changes to crops:

* There are eight trays ready to plant greens, The greens will be plant tomorrow.

* Some crops show better physiologycal conditions, maybe as a result of the drop of the temperature into the greeenhab.

Narrative:

The max temperature of the greenhab was reduced to 31°C, It’s an important reduction. However, it’s important try of reduce to 25°C or keep below to 30°C for avoid damage in the crops.

Rise the times for watering the crops from 3 to 5 times without change the volume of water, with the aim to icrease the moisture avariable on the crop’s soil along the day.

Watered Crew 201 experiments

Harvest: 7g Cilantro

Support/Supplies Needed:

Operations Report – January 14th

SOL: 2

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #9 still not working, Suits #3, 6 and 8 under charging review, a Suit charger is needed because there is one that does not work and no spares are available.

Notes on non-nominal systems: Suit #9 was repaired, the electrical connections were fixed and a broken 5A fuse was identified but it could not be changed because there is no spare fuses, everything but the fuse is operational in the suit at this moment. Suits #3, 6 and 8 are under charging review, pending for diagnostic.

A Suit charger and spare ones are needed.

Generator (hours run): 17hr 37min; Turned on last night (13Jan2019) at 17:40; Turned off this morning (14Jan2019) at 11:17 cloudy day…; Turned on tonight (14Jan2019) at 16:45

Solar SOC – Turned on (13Jan2019) 85%; Turned off (14Jan2019) 100%; Turned on (14Jan2019) 88%

Diesel Reading – Just under 50%

Propane Reading – Under 30%, almost 29%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 90%; 480 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 87%; 260 gallons

Water (loft) – At level marker 10

Static to Loft Pump used – Yes; At 8:45 to refill tank

Water Meter: 01400658

Toilet tank emptied: Yes, a few minutes ago and still feels the smell 😉

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used for training EVA´s 1 and 2

Hours: 67.7

Beginning charge: 100%

Ending charge: 90%

Currently charging: Yes

Opportunity rover used: Not used

Hours: 45.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes, used for training EVA’s 1 and 2

Hours: 69.7

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Refueled and returned by the Crew 202 with 172540 miles

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Suit #9 has a broken 5A fuse, still not working, Suits 3,6 and 8 under charging review, one charger is not working; Suit #7 has not been reviewed yet; I´m aware to make a diagnosis.

Summary of Hab operations: Some crew members after a hard review, based on their skills on different engineering areas, have identified that the gas-like smell is probably generated by the metal of the stairs, the welding residues and the oxide, we recommend to give a new layer of paint to the staircase to decrease the smell and protect the metal which is showing some signs of corrosion.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: We were informed this morning by Shannon that yesterday the power was tripped to the power stations because the rover plugs were not completely clean before they were plugged back into power, so the rovers were not charged, she fixed it this morning.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Low batteries on two smoke detectors, spare batteries are needed.

Questions, concerns and requests to Mission Support: Nothing to report

Astronomy Report – January 14th

Crew Astronomer: Santiago Vargas Dominguez

Description: Observing run aimed to complement information in Halpha for M31 data (newly processed image attached).

Some data acquired but observation is yet to be completed.

Observation keep active in the skynet observation list.

Objects Viewed: M31, Heart Nebula

Problems Encountered: none

Sol Summary – January 14th

Sol: 2

Summary Title: Settling in

Author’s name: Oscar Ojeda

Mission Status: Nominal

Sol Activity Summary: We’re getting comfortable in the station, we woke up, shared breakfast and talked about some details on the activity we would have today. Time flew as we had 2 EVAs scheduled, the first group left, made an exploration of the region close to campus, and returned to a hot lunch. Afterwards, the second EVA group left and repeated a similar exploration to group 1, consisting mainly of an on-suit performance assessment, and sample collection. The rest of the day was used to keep the station and rest a little.

Look Ahead Plan: Tomorrow the science begins…! the crew will stay at the hab to begin the work on the research, with some of the samples that were taken today.

Anomalies in work: None

Weather: Mostly cloudy (Cumulus), partially sunny at the afternoon. Temperature similar to yesterday.

Crew Physical Status: In good health, however, we’ve had 3 crewmembers with "more-than-normal" sneezing.

EVA: 2, Space suit performance asessment, and sample collection.

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report, EVA Report, Mission Plan, HSO Report.

Support Requested: We don’t have much salt left in the kitchen.